OpenStudy (anonymous):

The length of the base of a triangle is 8.5 cm and its height is 5.4 cm. If the area of the triangle is half of the product of its base length and height, the area of the triangle is __________ cm2.

OpenStudy (anonymous):

@Baseballa101 @hyntc

OpenStudy (anonymous):

The problem tells you that you find the area of triangles by finding ½ × base × height. The problem also tells you that for this triangle, the base is 8.5 cm and the height is 5.4 cm. So to find the area you multiply ½ × 8.5 × 5.4 That product is 22.95 square centimeters Answer: The area of the triangle is 22.95 cm²

OpenStudy (anonymous):

^

OpenStudy (studygurl14):

lenght of base = 8.5 height = 5.4 Area = 1/2(base times height) Area = 1/2(8.5 x 5.4)